### v2

# Practical and Accurate Local Edge Differentially Private Graph Algorithms

Abstract.
The rise of massive networks across diverse domains necessitates sophisticated graph analytics, often involving sensitive data and raising privacy concerns. This paper addresses these challenges using local differential privacy (LDP), which enforces privacy at the individual level, where no third-party entity is trusted, unlike centralized models that assume a trusted curator.
We introduce novel LDP algorithms for two fundamental graph statistics: $k$-core decomposition and triangle counting. Our approach leverages previously unexplored input-dependent private graph properties, specifically the degeneracy and maximum degree of the graph, to improve theoretical utility. Unlike prior methods, our error bounds are determined by the maximum degree rather than the total number of edges, resulting in significantly tighter theoretical guarantees. For triangle counting, we improve upon the previous work of Imola, Murakami, and Chaudhury (Imola et al., 2021, 2022), which bounds error in terms of the total number of edges. Instead, our algorithm achieves error bounds based on the graph’s degeneracy by leveraging a differentially private out-degree orientation, a refined variant of Eden et al.’s randomized response technique (Eden et al., 2023b), and a novel, intricate analysis, yielding improved theoretical guarantees over prior state-of-the-art.
Beyond theoretical improvements, we are the first to evaluate the practicality of local DP algorithms in a distributed simulation environment, unlike previous works that tested on a single processor. Our experiments on real-world datasets demonstrate substantial accuracy improvements, with our $k$-core decomposition achieving errors within $\mathbf{3}$x the exact values—far outperforming the $\mathbf{131}$x error in the baseline of Dhulipala et al. (Dhulipala et al., 2022) . Additionally, our triangle counting algorithm reduces multiplicative approximation errors by up to six orders of magnitude compared to prior methods, all while maintaining competitive runtime performance.
journalyear: 2024copyright: rightsretained
Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://doi.org/10.5281/zenodo.15741879.
1. Introduction
Figure 1. Local edge differential privacy (LEDP) Model
Graph statistics such as the $k$-core decomposition and triangle count provide important characteristics about the underlying graph, such as its well-connected communities. These analytics, often performed on sensitive and private graphs, are regularly shared with a wide audience, including researchers, companies, governments, and the public. As such, it is vital to investigate techniques that can safeguard these published graph statistics from privacy attacks.
The $k$-core decomposition assigns an “importance” value to each node, roughly representing its influence within the graph. It is widely used to analyze the structure of real-world graphs, including social, email, and disease transmission networks. Formally, a $k$-core of a graph is a maximal subgraph where the induced degree of every vertex in the subgraph is at least $k$. The $k$-core decomposition (see Definition 2.16 and Figure 2) assigns a number, denoted as $core(v)$, to each vertex $v$. This number, $core(v)$, represents the largest value of $k$ for which the $k$-core still includes vertex $v$. Unfortunately, these values pose privacy risks.
Consider the application of $k$-core decomposition to COVID transmission data (Qin et al., 2020; S Monteiro et al., 2021; Serafino et al., 2022; Guo et al., 2022; Chang and Ferrara, 2022) and other disease transmission networks (Ciaperoni et al., 2020) such as HIV (Fujimoto et al., 2022; Holme and Litvak, 2017). The core numbers are generated and published, sometimes even with location data (Serafino et al., 2020). Revealing the precise core numbers of every individual can lead to privacy breaches. Consider a scenario where exactly $c$ individuals have a core number of $c-1$. This implies they form a clique of $c$ vertices, all connected. In disease transmission graphs, this directly exposes a cluster of sensitive disease transmissions! Therefore, it’s essential to release privacy-preserving core numbers.
Similarly, triangle counting is widely used in applications that process sensitive data. The triangle count, which measures the number of three-node cycles in a graph, is a fundamental metric in community detection (Palla et al., 2005; Prat-Pérez et al., 2012; Sotiropoulos and Tsourakakis, 2021; Masuda et al., 2018), recommendation systems (Zeng et al., 2014; Liu et al., 2018), and clustering (Tseng et al., 2021). In databases, triangle counting is essential for graph analytics frameworks (Neo4j, 2012) and is leveraged in query optimization (Al Hasan and Dave, 2018; Bishnu et al., 2025a) and fraud detection (Sánchez-Aguayo et al., 2021). However, exposing triangle counts without privacy guarantees can lead to inference attacks that compromise user confidentiality. Recent works in security and privacy (Imola et al., 2022, 2021; Liu et al., 2024, 2022c, 2023a) have highlighted the risks associated with publishing triangle counts, reinforcing the importance of privacy-preserving graph analytics.
Approximate values for such statistics are widely used to improve scalability and efficiency with minimal utility impact. In graph databases and uncertain networks, approximate cores enable analysis under probabilistic edges (Bonchi et al., 2014); in social and recommendation systems, they support influence detection and improve accuracy (Kitsak et al., 2010; Ai et al., 2021). Often, approximate core numbers are used for preprocessing other algorithms such as clustering (Mei et al., 2021; Giatsidis et al., 2016). Triangle counts are used in clustering, fraud detection, and query optimization, where small errors are tolerable (Alon et al., 1997; Liu et al., 2023a; Bishnu et al., 2025b). In dynamic domains like cybersecurity and biology, approximations allow timely insights from noisy data (Srihari and Leong, 2013; Wang et al., 2025; Liu et al., 2023b). Approximate statistics with strong privacy guarantees are often practical and effective when these applications use sensitive data (Liu et al., 2023a; Dhulipala et al., 2022; Imola et al., 2021, 2022; Eden et al., 2023b; Liu et al., 2023b).
Differential privacy (DP) (Dwork et al., 2006) is often considered the “gold standard” in protecting individual privacy. Traditionally, DP has been studied in the central model, where a trusted curator has access to raw data and applies DP mechanisms before releasing the results. However, this assumption is often impractical, especially in modern systems that rely on decentralized or federated architectures. This motivates the local model of differential privacy, introduced by the seminal works of Evfimievski et al. (Evfimievski et al., 2003) and Kasiviswanathan et al. (Kasiviswanathan et al., 2011), which recently gained much attention in the theoretical computer science (Dhulipala et al., 2022; Eden et al., 2023b; Dinitz et al., 2024; Henzinger et al., 2024), cryptography and security (Imola et al., 2022, 2021; Fu et al., 2023; Li et al., 2023; Liu et al., 2022a; Wang et al., 2023b; Zhong et al., 2023; Liu et al., 2023a), data mining (Hillebrand et al., 2023a, b; Liu et al., 2024, 2022c; Murakami and Sei, 2023; Brito et al., 2023; Nguyen et al., 2024), query answering (Kulkarni, 2019; Tire and Gursoy, 2024; Wang et al., 2019; Farias et al., 2020; Roth et al., 2021), and machine learning (Wang et al., 2023a; Lin et al., 2022; Mueller et al., 2022; Hidano and Murakami, 2022; Zhu et al., 2023; Xiang et al., 2023; Seif et al., 2022; Guo et al., 2023; Hehir et al., 2022; Jiang et al., 2023) communities. In this model, each user independently applies DP mechanisms before sharing their privacy-preserving outputs with an untrusted curator. It offers stronger privacy guarantees by never exposing raw data and is naturally suited to distributed settings; it has been used in prominent cases including federated learning (Mahawaga Arachchige et al., 2022; Naseri et al., 2022; Kim et al., 2021), the 2020 U.S. Census (Abowd et al., 2022), and by companies such as Apple (Team, 2017).
While DP in the traditional database setting focuses on protecting individual records, many real-world datasets are inherently relational, represented as graphs. In these cases, the sensitive information are the edges, i.e., the connections between entities. This motivates the need for local differentially private (LDP) graph algorithms. The local edge differential privacy model (LEDP), as introduced in recent works (Qin et al., 2017; Imola et al., 2021; Dhulipala et al., 2022), represents a novel approach designed to ensure local privacy for graph outputs (Fig. 1). Graph data is increasingly integral to modern database systems, underpinning applications in knowledge graphs, social networks, cybersecurity, and financial fraud detection. Many relational databases now support graph extensions (e.g., SQL-based graph queries), while specialized graph databases (Neo4j, 2012; Bebee et al., 2018) are widely deployed in industry. However, applying LEDP algorithms in these settings is challenging: unlike in tabular data, where individual data points can be perturbed, perturbing edges within a graph introduces structural dependencies and high computational cost.
All previous implementations of local differentially private graph algorithms (Qin et al., 2017; Imola et al., 2023, 2022) use Randomized Response (RR) (Warner, 1965), which independently flips the presence or absence of each edge with a probability based on the privacy parameter $\varepsilon$. While simple and composable, RR introduces substantial noise, especially for small $\varepsilon\in(0,1]$, increasing the density [^1] The density is the ratio of the number of edges to the number of nodes. of the input graph and limiting both the accuracy and scalability of the algorithm. Dhulipala et al. (Dhulipala et al., 2022) proposed the first LEDP algorithm that goes beyond RR, leveraging the geometric mechanism for $k$-core decomposition. However, their algorithm is purely theoretical, and their error bounds scale poorly with graph size. In particular, they give additive error bounds $\geq\frac{\log_{2}^{3}(n)}{\varepsilon}$ (where $n$ is the number of vertices); on a graph with $n=10^{5}$ nodes and $\varepsilon=0.5$, this translates to an additive error of $9164$. Most real-world graphs of this size have max core numbers of $10^{2}$ magnitude, so the additive error itself leads to a $\geq 91$-multiplicative approximation factor—much too large for any practical use.
This work simultaneously develops both new theoretical and implementation techniques that, together with Randomized Response, enable provably private, accurate, and computationally efficient LEDP algorithms. We make the following contributions:
  • $\bullet$
    We design a novel LEDP $k$-core decomposition algorithm that doesn’t use Randomized Response and provides provable privacy and error guarantees. Leveraging the input-dependent maximum degree property of the graph, we achieve improved theoretical bounds over the LEDP $k$-core decomposition algorithm of Dhulipala et al. (Dhulipala et al., 2022) (see Table 1). Two key innovations lie in thresholding the maximum number of levels a node can move up based on its noisy (private) degree and the use of bias factors to reduce the impact of noise. Since a node’s core number is upper bounded by its degree, our algorithm offers stronger theoretical guarantees for most real-world graphs, where the maximum degree is significantly smaller than the number of nodes.
  • $\bullet$
    We present the first implementation of a private $k$-core decomposition algorithm and demonstrate through empirical evaluation that it achieves an average error of $\mathbf{3}$x the exact values, markedly improving upon prior approaches. Furthermore, our LEDP implementation attains error rates that closely align with the theoretical approximation bounds of the best non-private algorithms, underscoring its practical efficiency and accuracy.
  • $\bullet$
    We design a novel LEDP triangle counting algorithm that modifies our $k$-core decomposition to construct a low out-degree ordering, minimizing each node’s out-degree. Leveraging this ordering, our algorithm achieves improved theoretical error bounds over the best-known methods of Imola et al. (Imola et al., 2022, 2021) and Eden et al. (Eden et al., 2023b) for bounded degeneracy graphs, common in real-world networks (see Table 1). We present a novel analysis to analyze the non-trivial error bounds based on the Law of Total Expectation/Variance. Our implementation reduces relative error by up to 89x and improves the multiplicative approximation by up to six orders of magnitude over the best previous implementation (Imola et al., 2022), while maintaining competitive runtime.
  • $\bullet$
    Recognizing that the LEDP model (see Section 2.3) is inherently decentralized, we present the first evaluation of LEDP graph algorithms in a simulated distributed environment with actual message passing. Unlike prior studies that relied on a single processor, we simulate a distributed environment by partitioning the graph across multiple processors. This approach provides a more realistic assessment of both computational and communication overhead in large-scale distributed scenarios. We demonstrate the practicality of this evaluation by applying it to our $k$-core decomposition and triangle counting algorithms.
  • $\bullet$
    We present the first LEDP graph algorithm implementations that scale to billion-edge graphs, whereas prior implementations were tested on graphs with millions of edges (Imola et al., 2021, 2022). Our evaluation framework serves as a valuable tool for designing and testing other LEDP algorithms. Our source code is available at (git, 2024).
Previous Work This Work
$k$-Core $O\left(\frac{\log^{3}(n)}{\varepsilon}\right)$ (Dhulipala et al., 2022) $O\left(\frac{\log(D_{\max})\log^{2}(n)}{\epsilon}\right)$
Triangle Counting $O\left(\frac{n^{3/2}}{\varepsilon^{3}}+\frac{\sqrt{C_{4}}}{\varepsilon^{2}}\right)$ (Eden et al., 2023b) $O\left(\frac{\sqrt{n}d\log^{3}n}{\varepsilon^{2}}+\sqrt{\overrightarrow{C}_{4}}\right)$
Table 1. Additive error bounds compared to previous work. Here $n:$ number of nodes, $D_{max}:$ maximum degree, $\varepsilon:$ privacy parameter, $d:$ degeneracy, and $C_{4}$ and $\overrightarrow{C}_{4}$ is the number of $4$-cycles and oriented $4$-cycles, respectively.
1.1. Related Work
Local differential privacy (LDP) for graph data has been extensively studied (Qin et al., 2017; Imola et al., 2021, 2022; Sun et al., 2019; Ye et al., 2020b, a; Dhulipala et al., 2022; Eden et al., 2023b; Hillebrand et al., 2023a), focusing on tasks such as synthetic graph generation and subgraph counting. Some works (Sun et al., 2019; Liu et al., 2024) explore an extended local view, in which each node knows its full 2-hop neighborhood (i.e., its neighbors’ edges) to improve triangle counting accuracy. In that model, triangle counting is trivial since each node sees its entire set of incident triangles—unlike our model, where nodes see only immediate (one-hop) neighbors; hence, we require more complex algorithms since nodes cannot see their incident triangles in LEDP. Moreover, the extended view is often unrealistic, since users (e.g., in social networks) may not wish to reveal their private (potentially sensitive) friend lists to their friends.
The LEDP model was introduced by Qin et al. (Qin et al., 2017) and Imola, Murakami, and Chaudhury (Imola et al., 2021), with subsequent theoretical expansions (Dhulipala et al., 2022; Eden et al., 2023b; Imola et al., 2022). Imola et al. (Imola et al., 2021, 2022) provided the first practical LEDP implementations for triangle and subgraph counting. Recently, Hillebrand et al. (Hillebrand et al., 2023a) improved LEDP triangle counting using hash functions, though their method does not scale to large graphs. All prior LEDP triangle counting algorithms rely on Randomized Response. Imola et al. (Imola et al., 2021) introduced an LEDP triangle counting algorithm in both non-interactive (single-round) and interactive (multi-round) settings, bounding the standard deviation of the additive error by $O\left(\frac{n^{2}}{\varepsilon}+\frac{n^{3/2}}{\varepsilon^{2}}\right)$. In a subsequent work, they reduce the protocol’s communication cost (Imola et al., 2022) by using a combination of sampling and clipping techniques, and refined their standard deviation analysis by using the number of $4$-cycles, $C_{4}$. Their new theoretical standard deviation is $O\left(\frac{\sqrt{C_{4}}}{\varepsilon}+\frac{n^{3/2}}{\varepsilon^{2}}\right)$ for the interactive setting and $O(n^{2})$ for the non-interactive setting. Eden et al. (Eden et al., 2023b) further enhanced triangle counting accuracy with an improved post-processing analysis, achieving a standard deviation of $O\left(\frac{\sqrt{C_{4}}}{\varepsilon^{2}}+\frac{n^{3/2}}{\varepsilon^{3}}\right)$ for the non-interactive setting and establishing lower bounds of $\Omega(n^{2})$ and $\Omega\left(\frac{n^{3/2}}{\varepsilon}\right)$ for the non-interactive and interactive settings, respectively. Despite these advancements, prior work has neither combined Randomized Response with other privacy mechanisms to improve error bounds nor accounted for input-dependent properties of graphs in theoretical analyses.
For LEDP $k$-core decomposition, all known algorithms remain theoretical (Dhulipala et al., 2022). The algorithm by Dhulipala et al. (Dhulipala et al., 2022) uses a level data structure, where nodes ascend levels based on their noisy induced degrees, with noise drawn from a symmetric geometric distribution to ensure privacy. However, this noise scales with the number of nodes rather than adapting to input structure, leading to significant errors in large graphs. Recent concurrent and independent work by Dhulipala et al. (Dhulipala et al., 2024) introduces a generalized sparse vector technique to avoid cumulative privacy budget costs, achieving improved theoretical guarantees. However, implementing this approach in a distributed setting is challenging, as it relies on a peeling algorithm that is difficult to distribute. Additionally, the practical performance of these algorithms remains unexplored.
2. Preliminaries
Differential privacy on graphs is defined for edge-neighboring inputs. Edge-neighboring inputs are two graphs which differ in exactly one edge. Here, we consider undirected graphs.
Definition 2.1 (Edge-Neighboring (Nissim et al., 2007)).
Graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are edge-neighboring if they differ in one edge, namely, if $V_{1}=V_{2}$ and the size of the symmetric difference of $E_{1}$ and $E_{2}$ is 1.  [^2] The symmetric difference of two sets is the set of elements that are in either set, but not in their intersection.
With high probability (whp) is used in this paper to mean with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$.
The local edge differential privacy (LEDP) model assumes that each node in the input graph keeps their adjacency list private. The model is defined in terms of $\varepsilon$-DP algorithms, called $\varepsilon$-local randomizers ($\varepsilon$-LR), that are run individually by every node. The $\varepsilon$-LRs are guaranteed to be $\varepsilon$-DP when the neighboring inputs are adjacency lists that differ in one element. Following (Imola et al., 2022), we assume that the curator and all nodes act as honest-but-curious adversaries.
Definition 2.2 ($\varepsilon$-Edge Differential Privacy (Dwork et al., 2006; Nissim et al., 2007)).
[…]
$\frac{1}{e^{\varepsilon}}\leq\frac{\mathsf{Pr}[\mathcal{A}(G^{\prime})\in S]}{\mathsf{Pr}[\mathcal{A}(G)\in S]}\leq e^{\varepsilon}.$
Definition 2.3 (Local Randomizer (LR) (Dhulipala et al., 2022)).
An $\varepsilon$-local randomizer $R:\mathbf{a}\rightarrow\mathcal{Y}$ for node $v$ is an $\varepsilon$-edge DP algorithm that takes as input the set of its neighbors $N(v)$, represented by an adjacency list $\mathbf{a}=(b_{1},\dots,b_{|N(v)|})$. In other words,
$\frac{1}{e^{\varepsilon}}\leq\frac{\mathsf{Pr}\left[R(\mathbf{a}^{\prime})\in Y\right]}{\mathsf{Pr}\left[R(\mathbf{a})\in Y\right]}\leq e^{\varepsilon}$
for all $\mathbf{a}$ and $\mathbf{a}^{\prime}$ where the symmetric difference is $1$ and all sets of outputs $Y\subseteq\mathcal{Y}$. The probability is taken over the random choices of $R$ (but not over the choice of the input).
The previous definitions of LEDP (Qin et al., 2017; Imola et al., 2021; Dhulipala et al., 2022) are satisfied by the following Definition 2.4. (Dhulipala et al., 2022) gives a slightly more general and complex definition of LEDP in terms of transcripts but all of the algorithms in our paper satisfy our definition below, which is also guaranteed to satisfy their more general transcript-based definition.
Definition 2.4 (Local Edge Differential Privacy (LEDP) (Dhulipala et al., 2022)).
Given an input graph $G=(V,E)$, for any edge $\{u,v\}$, let algorithm $\mathcal{A}$ assign $\left((R^{u}_{1}(\mathbf{a}_{u},p_{1}),\varepsilon^{u}_{1}),\dots,(R^{u}_{r}(\mathbf{a}_{u},p_{r}),\varepsilon^{u}_{r})\right)$ to be the set of $\varepsilon^{u}_{i}$-local randomizers called by vertex $u$ during each interactive round and $\left((R^{v}_{1}(\mathbf{a}_{v},p_{1}),\varepsilon^{v}_{1}),\dots,(R^{v}_{s}(\mathbf{a}_{v},p_{s}),\varepsilon^{v}_{s})\right)$ be the set of $\varepsilon^{v}_{i}$-LRs called by $v$. The private adjacency lists of $u$ and $v$ are given by $\mathbf{a}_{u}$ and $\mathbf{a}_{v}$, respectively, and $p_{i}$ are the new public information released in each round. Algorithm $\mathcal{A}$ is $\varepsilon$-local edge differentially private (LEDP) if for every edge, $\{u,v\}$:
$\varepsilon^{u}_{1}+\cdots+\varepsilon^{u}_{r}+\varepsilon^{v}_{1}+\cdots+\varepsilon^{v}_{s}\leq\varepsilon.$
For intuition, each LR takes as input the private adjacency list of the node $v$ and public information released in previous rounds; then, it releases new public information for $v$ which will inform the computation of other nodes in the next round. Hence, the algorithm is interactive. Each time $v$ releases, it loses some amount of privacy indicated by $\varepsilon^{v}_{i}$ for the $i$-th LR. Since edge-neighboring graphs differ in exactly one edge, to ensure the privacy of the system, it is sufficient to ensure that the privacy loss of every edge sums up to $\varepsilon$. Thus, $\varepsilon$-LEDP algorithms also satisfy $\varepsilon$-DP (proven in (Dhulipala et al., 2022)).
2.1. Privacy Tools
We make use of the following privacy tools and primitives. We define all definitions below in terms of edge-neighboring adjacency lists since our tools will be applied to $\varepsilon$-local randomizers.
Definition 2.5 (Global Sensitivity (Dwork et al., 2006)).
For a function $f:\mathcal{D}\rightarrow\mathbb{R}^{d}$, where $\mathcal{D}$ is the domain of $f$ and $d$ is the dimension of the output, the $\ell_{1}$-sensitivity of $f$ is $GS_{f}=\max_{\mathbf{a},\mathbf{a}^{\prime}}\left\|f(\mathbf{a})-f(\mathbf{a}^{\prime})\right\|_{1}$ for all pairs of $\{\mathbf{a},\mathbf{a}^{\prime}\}$ of neighboring adjacency lists (differing in one neighbor).
Our algorithms and implementations in this paper use the symmetric geometric distribution defined in previous papers (Balcer and Vadhan, 2018; Chan et al., 2011; Dwork et al., 2006, 2010a; Farhadi et al., 2021; Shi et al., 2011). The symmetric geometric distribution is also often referred to as the “discrete Laplace distribution.” Using this distribution is quite crucial in practice in order to avoid the numerical errors associated with real-valued outputs from continuous distributions.
Definition 2.6 (Symmetric Geometric Distribution (Balcer and Vadhan, 2018; Shi et al., 2011)).
The symmetric geometric distribution, denoted $\mathsf{Geom}(b)$, with input parameter $b\in(0,1)$, takes integer values $i$ where the probability mass function at $i$ is given by $\frac{e^{b}-1}{e^{b}+1}\cdot e^{-|i|\cdot b}$.
We denote a random variable drawn from this distribution by $X\sim\mathsf{Geom}(b)$. With high probability (whp) is used in this paper to mean with probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$. As with all DP algorithms, privacy is always guaranteed and the approximation factors are guaranteed whp. We can upper bound the symmetric geometric noise whp using the following lemma.
Lemma 2.7 (Balcer and Vadhan, 2018; Chan et al., 2011; Dwork et al., 2006, 2010a; Farhadi et al., 2021; Shi et al., 2011; Dhulipala et al., 2022)).
With probability at least $1-\frac{1}{n^{c}}$ for any constant $c\geq 1$, we can upper bound $X\sim\mathsf{Geom}\left(x\right)$ by $|X|\leq\frac{c\ln n}{x}$.
The geometric mechanism is defined as follows.
Definition 2.8 (Geometric Mechanism (Balcer and Vadhan, 2018; Chan et al., 2011; Dwork et al., 2006, 2010a)).
Given any function $f:\mathcal{D}\rightarrow\mathbb{Z}^{d}$, where $\mathcal{D}$ is the domain of $f$ and $GS_{f}$ is the $\ell_{1}$-sensitivity of $f$, the geometric mechanism is defined as $\mathcal{M}(\mathbf{a},f(\cdot),\varepsilon)=f(\mathbf{a})+(Y_{1},\dots,Y_{d})$, where $Y_{i}\sim\mathsf{Geom}(\varepsilon/GS_{f})$ are independent and identically distributed (i.i.d.) random variables drawn from $\mathsf{Geom}(\varepsilon/GS_{f})$ and $\mathbf{a}$ is a private input adjacency list.
Definition 2.9 (Laplace Distribution).
The probability density function of the Laplace distribution on $X\in\mathbb{R}$ is $Lap(b)=2b\cdot\exp\left(-\left({\lvert X\rvert}\cdot b\right)\right)$
Lemma 2.10 (Laplace Mechanism (Dwork et al., 2006)).
Given a function $f:\mathcal{G}\rightarrow R$ with sensitivity $\Delta_{f}$, $f\left(\mathcal{G}\right)+Lap\left(\frac{\varepsilon}{\Delta_{f}}\right)$ is $\varepsilon$-differentially private.
Lemma 2.11 (Privacy of the Geometric Mechanism (Balcer and Vadhan, 2018; Chan et al., 2011; Dwork et al., 2006, 2010a)).
The geometric mechanism is $\varepsilon$-DP.
In addition to the Geometric Mechanism, our paper also uses Randomized Response (RR). Randomized Response (RR) when applied to graphs flips the bit indicating the existence of an edge in the graph. We define RR in terms of the way it is used on graphs.
Definition 2.12 (Randomized Response).
Randomized response on input graph $G=(V,E)$, represented as an upper triangular adjacency matrix $M$ which only contains entries $M[i,j]$ where $i<j$, flips every bit $M[i,j]$ (where $i<j$) in the matrix with probability $\frac{1}{e^{\varepsilon}+1}$.
It is well-known that randomized response is $\varepsilon$-DP.
Lemma 2.13 ((Dwork et al., 2006)).
Randomized response is $\varepsilon$-DP.
The composition theorem guarantees privacy for the composition of multiple algorithms with privacy guarantees of their own. In particular, this theorem covers the use case where multiple LEDP algorithms are used on the same dataset.
Theorem 2.14 (Composition Theorem (Dwork et al., 2006; Dwork and Lei, 2009; Dwork et al., 2010b)).
A sequence of DP algorithms, $(\mathcal{A}_{1},\dots,\mathcal{A}_{k})$, with privacy parameters $(\varepsilon_{1},\dots,\varepsilon_{k})$ form at worst an $\left(\varepsilon_{1}+\cdots+\varepsilon_{k}\right)$-DP algorithm under adaptive composition (where the adversary can adaptively select algorithms after seeing the output of previous algorithms).
Finally, the post-processing theorem states that the result of post-processing on the output of an $\varepsilon$-LEDP algorithm is $\varepsilon$-LEDP.
Theorem 2.15 (Post-Processing (Dwork et al., 2006; Bun and Steinke, 2016)).
Let $\mathcal{M}$ be an $\varepsilon$-LEDP algorithm and $h$ be an arbitrary (randomized) mapping from $Range(\mathcal{M})$ to an arbitrary set. The algorithm $h\circ\mathcal{M}$ is $\varepsilon$-LEDP.444$\circ$ is notation for applying $h$ on the outputs of $\mathcal{M}$.
We use implementations by the Google privacy team (Team, [n. d.]), which also guarantee cryptographic security.
2.2. Problem Definitions
Below, we define the $k$-core decomposition, low out-degree ordering, and triangle counting problems that we study.
Figure 2. Example $k$-core decomposition and triangles in a $4$-degenerate graph. Nodes are assigned core numbers based on the highest value core they belong to; e.g., a node in the $1$-core but not in the $2$-core is given the core number of $1$. Larger valued cores are contained within all smaller valued cores; e.g., the $3$-core is contained in the $1$ and $2$-core. Red edges show the triangles, i.e., $3$-cycles in the graph. The degeneracy of this graph is $4$.
In this paper, we consider undirected graphs $G=(V,E)$ with $n=|V|$ nodes and $m=|E|$ edges. We use $[n]$ to denote $\{1,\dots,n\}$. For ease of indexing, we set the IDs of $V$ to be $V=[n].$ The set of neighbors of a node $i\in[n]$ is denoted by $N(i)$, and the degree of node $i$ is denoted $\deg(i)$.
Definition 2.16 ($k$-Core Decomposition).
Given an input graph, $G=(V,E)$, a $k$-core is a maximal induced subgraph in $G$ where every node has degree at least $k$. A $k$-core decomposition assigns a core number to each node $v\in V$ equal to $\kappa$ if $v$ is in the $\kappa$-core but not the $(\kappa+1)$-core. Let $k(v)$ be the core number of $v$.
See Fig. 2 for an example. No exact $k$-core decomposition algorithm satisfies the definition of DP (or LEDP). Hence, our algorithms take an input graph $G$ and output an approximate core number for each node in the graph (Definition 2.17) and an approximate low out-degree ordering (Definition 2.19).
Definition 2.17 ($(\phi,\zeta)$-Approximate Core Number (Dhulipala et al., 2022)).
Let $\hat{k}(v)$ be an approximation of the core number of $v$, and let $\phi\geq 1,\zeta\geq 0$. The core estimate $\hat{k}(v)$ is a $(\phi,\zeta)$-approximate core number of $v$ if $k(v)-\zeta\leq\hat{k}(v)\leq\phi\cdot k(v)+\zeta$.
We define the related concept of an approximate low out-degree ordering based on the definition of degeneracy.
Definition 2.18 (Degeneracy).
An undirected graph $G=(V,E)$ is $d$-degenerate if every induced subgraph of $G$ has a node with degree at most $d$. The degeneracy of $G$ is the smallest value of $d$ for which $G$ is $d$-degenerate.
It is well known that degeneracy $d=\max_{v\in V}\{k(v)\}$.
Definition 2.19 ($(\phi,\zeta)$-Approximate Low Out-Degree Ordering).
Let $D=[v_{1},v_{2},\dots,v_{n}]$ be a total ordering of nodes in a graph $G=(V,E)$. The ordering $D$ is an $(\phi,\zeta)$-approximate low out-degree ordering if orienting edges from earlier nodes to later nodes in $D$ produces out-degree at most $\phi\cdot d+\zeta$.
Definition 2.20 (Triangle Count).
Given an undirected input graph $G=(V,E)$, the triangle count returns the number of $3$-cycles in $G$.
2.3. Distributed Simulation Model
Local Edge Differential Privacy (LEDP) is inherently decentralized: each user (or node) independently perturbs their private local data (adjacency list) before any communication. This model aligns naturally with distributed systems, where data is often siloed across machines or clients. To evaluate LEDP algorithms in such settings, we adopt a distributed simulation framework that closely mirrors real-world deployments. Specifically, we use a coordinator-worker model in which each worker is assigned a partition of nodes along with their full adjacency lists. Workers execute LEDP algorithms locally and communicate only privacy-preserving outputs to a central coordinator. The coordinator aggregates these responses and broadcasts public updates to all workers, proceeding iteratively over synchronous communication rounds. While assigning one machine per node is infeasible at scale, this simulation preserves the privacy and communication structure of LEDP and allows for practical evaluation of network overhead on large graphs.
3. Practical $k$-Core Decomposition Algorithm
We present $k$-CoreD, a novel $k$-core decomposition algorithm that addresses key limitations of prior work (Dhulipala et al., 2022) through principled algorithmic design. While their framework offers strong theoretical foundations under the $\epsilon$-LEDP model, its dependence on the total number of nodes leads to large additive error and excessive communication rounds. Our algorithm replaces this dependency with input-sensitive parameters—specifically, the graph’s maximum degree—through degree thresholding and bias terms. These techniques yield improved asymptotic bounds and significantly lower empirical error, as confirmed by our experiments.
Figure 3. Node movements in $k$-CoreD’s Level Data Structure (LDS). Green: active nodes eligible to move; red: thresholded nodes; orange: active nodes that fail the noisy neighbor check. The LDS and threshold structures are shown alongside the graph. Noise is added during the level-moving step to ensure privacy, and snapshots illustrate node progression and halted movement due to thresholding.
3.1. Algorithm Description
Our algorithm operates synchronously over $O\left(\log(n)\log\left(D_{\max}\right)\right)$ rounds, where $D_{\max}$ is the maximum degree of the graph. The algorithm outputs $\left(2+\eta,O\left(\frac{\log(D_{\max})\log^{2}(n)}{\epsilon}\right)\right)$-approximate core numbers with high probability, as well as a low out-degree ordering with the same approximation guarantee. Throughout this section, the term $\log(n)$ denotes $\log_{1+\psi}(n)$, unless explicitly stated otherwise (where $\psi$ is a constant). The algorithm uses a level data structure (LDS) (Dhulipala et al., 2022), where nodes are assigned levels that are updated iteratively. Levels are partitioned into groups of equal size, with each group $g_{i}$ containing $\frac{\lceil\log(n)\rceil}{4}$ consecutive levels. We limit the number of rounds a node participates in, based on its noisy degree, which we refer to as degree thresholding. This significantly reduces the number of rounds, from $O(\log^{2}(n))$ (Dhulipala et al., 2022) to $O\left(\log(n)\log\left(D_{\max}\right)\right)$. In each round, the algorithm uses a noisy count of a node’s neighbors at the same level to decide if it should move up a level. After processing all nodes in a round, the updated LDS is published for use in subsequent rounds. Once all rounds are complete, the algorithm estimates the core numbers of the nodes based on their final levels, using Algorithm 4. Additionally, a low out-degree ordering is determined by sorting nodes from smaller to larger levels, breaking ties using node IDs. [^5] This doesn’t leak privacy, as IDs are assigned to nodes and not edges and reveal no information about the sensitive edge data.. The algorithm is implemented in a distributed setting, where computation is divided between a coordinator and multiple workers. The pseudocode is structured to reflect this division. We now describe their respective roles.
Coordinator. As described in Algorithm 1, the coordinator takes as input the graph size $n$, number of workers $M$, constant $\psi>0$, privacy parameter $\varepsilon\in(0,1]$, privacy split fraction $f\in(0,1)$, and bias term $b$. It first computes the privacy budgets $\varepsilon_{1}$ and $\varepsilon_{2}$ for degree thresholding and noisy neighbor counts, respectively, based on $\varepsilon$ and $f$ (Algorithm 1). The coordinator maintains the level data structure (LDS), where $LDS[i]$ stores the current level of node $i$, and a communication channel, channel, for receiving messages from the workers. All nodes are initialized to level 0 (Algorithm 1) and are incrementally moved up in later rounds based on signals received from the workers. It begins by signaling the workers to load their assigned subgraphs in parallel (Algorithm 1) and then collects the degree threshold values to determine the total number of rounds (Algorithm 1). In each round $r$, it computes the corresponding group index (Algorithm 1) and launches $M$ asynchronous worker processes (Algorithm 1). Each worker returns a bit vector indicating whether each node in its subgraph should move up a level. After all processes complete (Algorithm 1), the coordinator processes the responses and updates the LDS accordingly. It then publishes the new LDS (Algorithm 1) before the next round begins. After the final round, the coordinator computes the estimated core numbers using the final LDS.
Worker (Degree Thresholding). As shown in Algorithm 2, each worker begins by loading its assigned subgraph and initializing local structures. For each node $v$, it computes a noisy degree $\widetilde{d}_{v}=d_{v}+X$, where $d_{v}$ is the true degree and $X\sim\text{Geom}(\frac{\epsilon}{2})$ is symmetric geometric noise (Algorithm 2). To mitigate large positive noise and reduce overestimation, a bias term—proportional to the noise’s standard deviation—is subtracted from $\widetilde{d}_{v}$. The worker then computes a threshold value for each node as $\left\lceil\log_{2}(\widetilde{d}_{v})\right\rceil\cdot L$, where $L$ is the number of levels per group in the LDS (Algorithm 2). These thresholds determine the maximum number of rounds in which each node participates. The worker keeps track of the maximum threshold across its subgraph and returns it to the coordinator (Algorithm 2).
Worker (Level Moving). In each round, workers assess whether nodes in their subgraph should move up a level. As shown in Algorithm 3, if a node $v$ has already reached its threshold round $r$, it is skipped (Algorithm 3). Otherwise, if $v$ is currently at level $r$, we count the number of its neighbors that are also at the same level (Algorithm 3). To ensure privacy, this count $\mathcal{U}_{v}$ is perturbed to produce a noisy estimate $\widehat{\mathcal{U}}_{v}=\mathcal{U}_{v}+X+B$ (Algorithm 3), where $X$ is symmetric geometric noise with parameter $s=\frac{\varepsilon}{2\cdot v.\textit{threshold}}$ and $B$ is an added bias term to counteract large negative noise. The node moves up a level if $\widehat{\mathcal{U}}_{v}>(1+\eta/5)^{\mathcal{F}(r)}$, where $\mathcal{F}(r)$ is the group index corresponding to the current level (Algorithm 3). After processing all nodes, the worker sends the updated level-change bits to the coordinator (Algorithm 3).
Bias Terms. We introduce two analytically derived bias terms, based on the standard deviation of the symmetric geometric distribution—one for degree thresholding and one for level movement. The first bias term is subtracted from the computed threshold to account for situations where a large positive noise is chosen. If a large positive noise is chosen, we lose privacy proportion to the new threshold in the level moving step. Hence, our bias term biases the result to smaller thresholds, resulting in less privacy loss.
The second bias term is added to the computed induced noisy degree to account for situations where a large negative noise prevents nodes from moving up the first few levels of the structure. Our added bias allows nodes with non-zero degrees to move up the first levels of the structure. Since the degree bounds increase exponentially, this additional bias term accounts for smaller errors as nodes move up levels. Notably, we observe this behavior in our experiments when comparing to the baseline implementation of (Dhulipala et al., 2022), which omits the bias term: many nodes remain stuck at level 0, resulting in significantly higher error. This highlights the practical importance of our bias correction.
Example 3.1.
Fig. 3 illustrates node movement in the LDS during $k$-CoreD. In each round, nodes compute a noisy count of same-level neighbors and move up if it exceeds a threshold based on group index—unless blocked by their degree threshold. In the LDS, green marks eligible nodes; red in the threshold array marks those blocked; and in the graph view, green means movement and red/orange means restriction. For instance, node $c$ is blocked in Round 1, $a$ in Round 2, and $b$ in Round 3. In Round 2, node $d$ is not thresholded but remains at the same level due to failing the noisy neighbor check (Algorithm 3 Algorithm 3).
1 Input: graph size $n$; number of workers $M$; approx constant $\psi\in\left(0,1\right)$; privacy parameter $\varepsilon\in(0,1]$; split fraction $f\in(0,1)$; bias term $b$.
2 Output: Approximate core numbers and low out-degree ordering of each node in $G$.
3 Function $k$-CoreD($n,\psi,\varepsilon,f,b$)
4       Set $\lambda=\frac{(5-2\eta)\eta}{(\eta+5)^{2}};L=\frac{\left\lceil{\log n}\right\rceil}{4}$
5       Set $\varepsilon_{1}=f\cdot\varepsilon$ and $\varepsilon_{2}=\left(1-f\right)\cdot\varepsilon$
6       Set $\mathcal{C}\leftarrow$ new $\text{Coordinator}\left(LDS,channel\right)$
7       Coordinator initializes $\mathcal{C}.LDS$ with $\mathcal{C}.LDS[i]\leftarrow 0\ \forall i\in[n]$.
[…]
9       parfor $w=1$ to $M$ do
10             maxDegreeThresholds$\left[w\right]=\textnormal{{DegreeThresholdWorker}}(w,\varepsilon_{1},L,b)$
11
[…]
13       for $r=0$ to numOfRounds do
14             Set $\mathcal{F}(r)\leftarrow\left\lfloor\frac{r}{L}\right\rfloor$
15             parfor $w=1$ to $M$ do
[…]
17
18            $\mathcal{C}.\text{wait}()$triangleright coordinator waits for workers to finish
19             nextLevels $\leftarrow\mathcal{C}.channel$
20             for $i=1$ to $n$ do
21                   if nextLevels$\left[i\right]=1$ then
22                         $\mathcal{C}.LDS.\text{levelIncrease}\left(i\right)$
23
24
25            Coordinator publishes updated $\mathcal{C}.LDS$
26
27      Coordinator calls $cores\leftarrow\mathcal{C}.\textnormal{{EstimateCoreNumbers}}(\mathcal{C}.LDS,L,\lambda,\psi)$
28       Coordinator produces $D$, a total order of all nodes, using levels from $\mathcal{C}.LDS$ (from smaller to larger) breaking ties by node ID
29       Return $(cores,D)$
Algorithm 1 $k$-Core Decomposition and Ordering (Coordinator)
1 Input: worker ID $w$; privacy parameter $\varepsilon\in(0,1]$; levels per group $L$; bias term $b$.
2 Function DegreeThresholdWorker($w,\varepsilon,L,b$)
3       Set maxThreshold0absent0$\leftarrow$ 0← 0
4       for node ${}v:=$ localGraph do
5             Sample $X\sim\mathsf{Geom}\left(\frac{\varepsilon}{2}\right)$
6             $\widetilde{d}_{v}\leftarrow d_{v}+X$triangleright noised degree
7             $\widetilde{d}_{v}\leftarrow\widetilde{d}_{v}+1-\min\left(b\cdot\frac{2\cdot e^{\varepsilon}}{e^{2\varepsilon}-1},\widetilde{d}_{v}\right)$
8             $v$.threshold $\leftarrow\left\lceil{\log_{2}(\widetilde{d}_{v})}\right\rceil\cdot L$triangleright thresholding
9             $v$.permZero $\leftarrow 1$
[…]
11
12      $w.\text{send}\left(\text{\emph{maxThreshold}}\right)$
13
Algorithm 2 Degree Thresholding (Worker)
1 Input: worker id $w$; round number $r$; privacy parameter $\varepsilon\in(0,1]$; constant $\psi$; group index $\mathcal{F}(r)$; pointer to the coordinator $LDS$.
2 Function LevelMovingWorker($w,r,\varepsilon,\psi,\mathcal{F}(r),LDS$)
3       Set nextLevels $\leftarrow[0,\ldots,0]$
4       for node ${}v:=$ localGraph do
5             if v.threshold $=r$ then
6                   $v$.permZero = 00
7
8
9            vLevel $\leftarrow LDS.\text{getLevel}\left(v\right)$
10             if vLevel $=r$ and v.permZero $\neq 0$ then
11                   Let $\mathcal{U}_{v}$ be the number of neighbors $j\in\mathbf{a}_{v}$ where $LDS.\text{getLevel}\left(j\right)=r$.
12                   Set scale $s\leftarrow\frac{\varepsilon}{2\cdot\left(v.\text{threshold}\right)}$
13                   Sample $X\sim\mathsf{Geom}(s)$.
14                   Set extra bias $B\leftarrow\frac{6e^{s}}{\left(e^{2s}-1\right)^{3}}$
15                   Compute $\widehat{\mathcal{U}}_{v}\leftarrow\mathcal{U}_{v}+X+B$.
16                   if $\widehat{\mathcal{U}}_{i}>(1+\eta/5)^{\mathcal{F}(r)}$ then
17                         nextLevels$\left[v\right]=1$
18
19                  else
20                         $v$.permZero $=0$
21
22
23
24      $w.\text{send}\left(w,\text{\emph{nextLevels}}\right)$
25       $w.\text{done}\left(\right)$
Algorithm 3 Level Moving (Worker)
1 Function EstimateCoreNumbers($LDS,L,\lambda,\eta$)
2       for $i=1$ to $n$ do
3             $\hat{k}(i)\leftarrow(2+\lambda)(1+\eta/5)^{\max\left(\left\lfloor\frac{LDS[i]+1}{L}\right\rfloor-1,0\right)}$.
4      Return $\{(i,\hat{k}(i)):i\in[n]\}$
Algorithm 4 Estimate Core Number (Coordinator) (Liu et al., 2022b)
3.2. Theoretical Analysis
Memory Analysis & Communication Cost
Let $M$ be the number of workers and $n$ the graph size. Each worker processes $S$ nodes, where $S=\lfloor n/M\rfloor$ for $M-1$ workers, and the last worker handles $n-(M-1)\lfloor n/M\rfloor$. The coordinator maintains the level data structure (LDS) and a communication channel, both requiring $O(n)$ space, resulting in a total memory usage of $O(n)$. Each worker processes $O(S)$ nodes, requiring $O(Sn)$ space for the graph and an additional $O(S)$ space for auxiliary structures, leading to a total of $O(Sn)$. In terms of communication, workers send one bit per node per round, incurring a per-worker cost of $O(S)$ and an overall round cost of $O(n)$. The coordinator receives and distributes the updated LDS, adding another $O(n)$ cost. Thus, the total communication overhead for the algorithm is $O\left(n\log(n)\log\left(D_{\max}\right)\right)$.
Privacy Guarantees
Our privacy guarantees depend on the following procedures. First, we perform degree-based thresholding, which upper bounds the number of levels a node can move up. Second, we subtract and add bias terms to the results of our mechanisms. And finally, we scale our noise added in Algorithm 3 of Algorithm 3 by the noisy threshold. We show that our algorithm can be implemented using local randomizers (Definition 2.3). Then, we show that the local randomizers have appropriate privacy parameters to satisfy $\varepsilon$-LEDP (Definition 2.4).
Lemma 3.2 (Degree Threshold LR).
Our degree thresholding procedure run with privacy parameter $\varepsilon^{\prime}$ is a $(\varepsilon^{\prime}/2)$-local randomizer.
Proof.
Our degree-thresholding procedure upper bounds the number of levels that we iterate through using the (private) degree of each node. Specifically, it adds symmetric geometric noise to the degree $\widetilde{d}_{u}=d_{u}+\mathsf{Geom}(\varepsilon^{\prime}/2)$ and then computes $\lceil\log_{1+\eta}(\widetilde{d}_{u})\rceil\cdot L$, where $L$ is the number of levels per group. The sensitivity of the degree of any node is $1$ and by the privacy of the geometric mechanism (Lemma 2.11), the output $\widetilde{d}_{u}$ is $(\varepsilon^{\prime}/2)$-DP. Then, producing the final level upper bound uses post-processing (Theorem 2.15) where privacy is preserved. Hence, our output is $(\varepsilon^{\prime}/2)$-DP and the algorithm can be implemented as a $(\varepsilon^{\prime}/2)$-local randomizer. ∎
Using Lemma 3.2, we prove Theorem 3.3.
Theorem 3.3.
Algorithm 1 is $\varepsilon$-LEDP.
Proof.
Our algorithm calls the local randomizers in Lemma 3.2 with $\varepsilon_{1}=\varepsilon\cdot f$, where $f\in(0,1)$ is a fraction which splits some portion of the privacy budget, and then iterates through the levels one-by-one while adding noise to the induced degree of each node consisting of all neighbors of the node on the same or higher level. We showed in Lemma 3.2 that the degree thresholding procedure can be implemented as $(\varepsilon_{1}/2)$-local randomizers.
The key to our better error bounds is our upper bound on the number of levels we iterate through, bounded by our threshold. Since the thresholds are public outputs from the local randomizers, we can condition on these outputs. Let the threshold picked for node $v$ be denoted as $t_{v}$. Then, we add symmetric geometric noise to the induced degree of the node (among the neighbors at or above $v$’s current level) drawn from $\mathsf{Geom}\left(\varepsilon_{2}/(2\cdot t_{v})\right)$ where $\varepsilon_{2}=\varepsilon\cdot\left(1-f\right)$. Conditioning on the public levels of each node, the sensitivity of the induced degree of any node is $1$. By the privacy of the geometric mechanism, we obtain a $\left(\varepsilon_{2}/(2\cdot t_{v})\right)$-local randomizer for $v$. By composition (Theorem 2.14) over at most $t_{v}$ levels, the set of all local randomizers called on $v$, is $(\varepsilon_{2}/2)$-differentially private. For any edge, the sum of the privacy parameters of the set of all local randomizers called on the endpoints of the edge is $2\cdot\varepsilon_{1}/2+2\cdot\varepsilon_{2}/2=\varepsilon_{1}+\varepsilon_{2}=f\cdot\varepsilon+(1-f)\cdot\varepsilon$. By Definition 2.4, this is $\varepsilon$-LEDP.
Finally, our bias terms, added or subtracted after applying the geometric mechanism, preserve privacy due to the post-processing invariance of differential privacy (Theorem 2.15). ∎
Approximation Guarantees
Our algorithm given in the previous section contains several changes that results in better theoretical bounds and optimizes the practical performance on real-world datasets. To prove our approximation factors, we first show Invariant 1 and Invariant 2 hold for our modified algorithm. $D_{\max}$ is the graph’s max degree.
Invariant 1 (Degree Upper Bound (Dhulipala et al., 2022)).
If node $i\in V_{r}$ (where $V_{r}$ contains nodes in level $r$) and $r<4\log^{2}n-1$, then $i$ has at most $(1+\eta/5)^{\lfloor r/(2\log n)\rfloor}+\frac{c\log(D_{\max})\log^{2}(n)}{\varepsilon}$ neighbors in levels $\geq r$, with high probability, for constant $c>0$.
Invariant 2 (Degree Lower Bound (Dhulipala et al., 2022)).
If node $i\in V_{r}$ (where $V_{r}$ contains nodes in level $r$) and $r>0$, then $i$ has at least $(1+\eta/5)^{\lfloor(r-1)/(2\log n)\rfloor}-\frac{c\log(D_{\max})\log^{2}(n)}{\varepsilon}$ neighbors in levels $\geq r-1$, with high probability, for constant $c>0$.
There are two parts to our analysis: first, we prove that our degree thresholding procedure does not keep the node at too low of a level, with high probability; second, we show that our new procedure for moving up levels adds at most the noise used in (Dhulipala et al., 2022) and not more. Together, these two arguments maintain the invariants.
Lemma 3.4.
Degree-thresholding satisfies Invariant 1.
Proof.
By Lemma 2.7, the noise we obtain for thresholding is upper bounded by $\frac{c^{\prime}\ln n}{(\varepsilon_{1}/2)}=\frac{2c^{\prime}\ln n}{f\cdot\varepsilon}$ with probability at least $1-\frac{1}{n^{c^{\prime}}}$. Thus, the noisy degree $\tilde{d_{v}}$ we obtain in Algorithm 2 of Algorithm 2 follows $\tilde{d_{v}}\geq d_{v}-\frac{2c^{\prime}\ln n}{f\cdot\varepsilon}$ with probability at least $1-\frac{1}{n^{c^{\prime}}}$. Let $r$ be the level we output as the threshold level. Then, we can compute the upper degree bound of this level to be at least $(1+\eta/5)^{\lfloor r/(2\log n)\rfloor}=(1+\eta/5)^{\log_{1+\eta/5}(\tilde{d_{v}})}\geq(1+\eta/5)^{\log_{1+\eta/5}\left(d_{v}-\left(\frac{2c^{\prime}\ln n}{f\varepsilon}\right)\right)}=d_{v}-\frac{2c^{\prime}\ln n}{f\varepsilon}$. The maximum induced degree of $v$ on any level is at most $d_{v}$. Let $d_{v,r}$ be the induced degree of $v$ on the thresholded level $r$, it must hold that $d_{v,r}\leq\left(d_{v}-\frac{2c^{\prime}\ln n}{f\varepsilon}\right)+\frac{4c^{\prime}\ln n}{f\varepsilon}+\frac{2c_{3}\ln n}{f\varepsilon}\leq(1+\eta/5)^{\lfloor r/(2\log n)\rfloor}+\frac{(4c^{\prime}+2c_{3})\ln n}{f\varepsilon}$. Since $f$ and $c^{\prime}$ are both constants and Invariant 1 allows for picking a large enough constant $c>0$, we can pick $c\geq 2(c^{\prime}+c_{3})/f$ and Invariant 1 is satisfied where $c^{\prime}$ and $c_{3}$ are fixed constants $\geq 1$. ∎
We do not have to prove that our thresholded level satisfies Invariant 2 since we use the threshold level as an upper bound of the maximum level that a node can be on. Hence, a node will not reach that level unless the procedure for moving the node up the levels satisfies Invariant 2. We now prove that our level movement procedure satisfies both invariants.
Lemma 3.5.
Our level moving algorithm satisfies Invariant 1 and Invariant 2.
Proof.
Our level moving algorithm is similar to (Dhulipala et al., 2022) except that we pick noise based on the threshold. Thus, by Lemma 2.7, our algorithm picks noise that is at most $\frac{2c_{1}\cdot t_{v}\ln n}{\varepsilon}$ with probability at least $1-\frac{1}{n^{c_{1}}}$ where $t_{v}$ is the released threshold for $v$ and $c_{1}\geq 1$ is a fixed constant. Also, by Lemma 2.7, it holds that $t_{v}\leq\log\left(D_{\max}+\frac{2c_{2}\ln n}{f\varepsilon}-\frac{2c_{3}\ln n}{f\varepsilon}\right)\cdot L\leq 2\log\left(D_{\max}\right)\log n$ with probability at least $1-\frac{1}{n^{c_{2}}}$. A node moves up a level from level $r$ if its induced degree plus the noise exceeds the threshold $(1+\eta/5)^{\lfloor r/(2\log n)\rfloor}$. Thus, using our computed noise, the degree of a node must be at least $(1+\eta/5)^{\lfloor r/(2\log n)\rfloor}-2\log\left(D_{\max}\right)\left(\frac{2c_{1}\log^{2}n}{\varepsilon}\right)$ with probability at least $1-\frac{1}{n^{c_{1}}}-\frac{1}{n^{c_{2}}}$ when it moves up from level $r$. By choosing large enough constants $c_{1},c_{2}>0$ and $c\geq 4c_{1}c_{2}$, we satisfy Invariant 2. Similarly, the node does not move up from level $r$ when its induced degree plus noise is at most $(1+\eta/5)^{\lfloor r/(2\log n)\rfloor}$. By a symmetric argument to the above, we show that Invariant 1 is satisfied. ∎
Finally, using Lemma 3.4 and Lemma 3.5, we can prove the final approximation factor for our algorithm.
Theorem 3.6.
Our algorithm returns $(2+\eta,O(\log(D_{\max})\log^{2}(n)/\varepsilon))$-approximate $k$-core numbers, with high probability, in $O\left(\log(n)\log\left(D_{\max}\right)\right)$ rounds of communication.
Proof.
By Lemma 3.4 and Lemma 3.5, our algorithm satisfies Invariant 1 and Invariant 2, hence, our algorithm returns our desired approximation using Theorem 4.1 of (Dhulipala et al., 2022). Our algorithm iterates to value at most $O(\log n\log\left(D_{\max}\right))$ for $D_{\max}$ (the maximum threshold), resulting in a total of $O\left(\log n\log\left(D_{\max}\right)\right)$ rounds. ∎
4. Triangle Counting using Low Out-Degree Ordering
We present our novel triangle counting algorithm, $\text{EdgeOrient}_{\Delta}$, which leverages the low out-degree ordering obtained from the $k$-CoreD algorithm along with randomized response (RR). Prior works rely on either all or a sample of neighboring edges after applying RR, and often suffer from error bounds that scale poorly with graph size. To address these limitations, our algorithm exploits a new input-dependent graph property, the degeneracy (maximum core number), to upper bound the number of oriented 4-cycles (Fig. 5), yielding significantly tighter error bounds in theory and practice.
1 Input: graph size $n$; number of workers $M$; constant $\psi\in\left(0,1\right)$; privacy parameter $\varepsilon\in(0,1]$; split fraction $f\in(0,1)$; bias term $b$;
2 Output: Noisy Triangle Count
3 Function $\text{EdgeOrient}_{\Delta}$($n,M,\psi,\varepsilon,f,b$)
4       Set $Z\leftarrow\textnormal{{$k$-CoreD}}(n,\psi,\frac{\varepsilon}{4},f,b)$ (Algorithm 1)
5       Set $\mathcal{C}\leftarrow$ new $\text{Coordinator}\left(cRR,cTCount,cMaxOut,X\right)$
6       $\triangleright$ Round 1: Randomized Response
7       parfor $w=1$ to $M$ do
8             $\textnormal{{RRWorker}}\left(w,n,\frac{\varepsilon}{4}\right)$
9
10      $\mathcal{C}.\text{wait}()$triangleright coordinator waits for workers to finish
11       $\mathcal{C}.\text{publishNoisyEdges}(cRR,X)$
12       $\triangleright$ Round 2: Max Out-degree
13       parfor $w=1$ to $M$ do
[…]
15
16      $\mathcal{C}.\text{wait}()$
17       $\widetilde{d}_{\max}\leftarrow\max\left(\{C.cMaxOut[i]\mid i\in[M]\}\right)+\frac{12\log(n)}{\varepsilon}$
18       $\triangleright$ Round 3: Count Triangles
19       parfor $w=1$ to $M$ do
[…]
21
22      $\mathcal{C}.\text{wait}()$
23       $\widetilde{\Delta}\leftarrow\sum_{i=1}^{n}C.cTCount\left[i\right]$
24       Return $\widetilde{\Delta}$
Algorithm 5 $\varepsilon$-LEDP Triangle Counting (Coordinator)
1 Input: worker id $w$; graph size $n$; privacy parameter $\varepsilon\in(0,1]$;
2 Function RRWorker($w,n,\varepsilon$)
3       Set neighborsRR [][$\leftarrow$[][]← [ ] [ ]
4       for node ${}v:=$ localGraph do
5             For all $ngh_{v}\leftarrow\left\{j:j\in\left[n\right]\wedge j>v\right\}$
6             neighborsRR$\left[v\right]=\text{RandomizedResponse}_{\varepsilon}(ngh_{v})$
7
8      $w.\text{sendRR}(w,\text{\emph{neighborsRR}})$
9       $w.\text{done}\left(\right)$
Algorithm 6 Randomized Response (Worker)
1 Input: worker id $w$; graph size $n$; privacy parameter $\varepsilon\in(0,1]$;
2 Function MaxOutDegreeWorker($w,\varepsilon,Z$)
3       Set $out_{\max}\leftarrow 0$.
4       for node, adjacency list ${}v,ngh:=$ localGraph do
5             Set $ngh_{v}\leftarrow\left\{j:j\in ngh\wedge Z[j]>Z[v]\right\}$
6             $B\sim\mathsf{Geom}(\varepsilon)$
7             $out_{\max}\leftarrow\max(out_{\max},|ngh_{v}|+B)$
8
9      $w.\text{sendMaxOutdegree}(out_{\max})$
10       $w.\text{done}\left(\right)$
Algorithm 7 Noisy Max Out-Degree (Worker)
1 Input: worker id $w$; privacy parameter $\varepsilon\in(0,1]$; low out-degree ordering $Z$, published noisy edges $X$; public maximum noisy out-degree $\widetilde{d}_{\max}$;
2 Function CountTrianglesWorker($w,\varepsilon,Z,X,\widetilde{d}_{\max}$)
3       Set workerTCount $\leftarrow 0.0$
4       for node, adjacency list ${}v,ngh:=$ localGraph do
5             Set $\widetilde{\Delta}\leftarrow 0.0$
6             $\text{OutEdges}_{v}=\left\{j:j\in ngh\wedge Z\left[j\right]>Z\left[v\right]\right\}$
7             for $i_{1}\in\{1,\dots,\min(\widetilde{d}_{\max},|\text{OutEdges}_{v}|)\}$ do
8                   for $i_{2}\in\{i_{1}+1,\dots,\min(\widetilde{d}_{\max},|\text{OutEdges}_{v}|)\}$ do
9                         $j\leftarrow\text{OutEdges}_{v}[i_{1}]$
10                         $k\leftarrow\text{OutEdges}_{v}[i_{2}]$
11                         $\widetilde{\Delta}\leftarrow\widetilde{\Delta}+\frac{X_{\left\{j,k\right\}}\cdot\left(e^{\varepsilon}+1\right)-1}{e^{\varepsilon}-1}$
12
13
14            Sample $R\sim\text{Lap}\left(\frac{\varepsilon}{2\cdot\widetilde{d}_{\max}}\right)$
15             $\widetilde{\Delta}\leftarrow\widetilde{\Delta}+R$
16             workerTCount $\leftarrow workerTCount+\widetilde{\Delta}$
17      $w.\text{sendTCount}\left(w,\text{\emph{workerTCount}}\right)$
18       $w.\text{done}\left(\right)$
Algorithm 8 Triangle Counting (Worker)
Figure 4. $\text{EdgeOrient}_{\Delta}$ for blue node: true triangle count is $2$, but due to Randomized Response and low out-degree ordering, estimate is $1$.
4.1. Algorithm Description
Our algorithm consists of three additional computation rounds after computing the low out-degree ordering ($Z$) using $k$-CoreD. In the first round, each node perturbs its adjacency list using Randomized Response (RR) (Warner, 1965), producing a privacy-preserving set of noisy edges for subsequent computations. In the second round, we calculate the maximum noisy out-degree, $\widetilde{d}_{\max}$, by determining each node’s outgoing edges based on $Z$. While the first two rounds can be combined, we separate them for clarity. In the final round, we compute the number of triangles incident to each node, using the noisy edges and the maximum noisy out-degree, $\widetilde{d}_{\max}$. The algorithm is implemented in a distributed setting, where computation is divided between a coordinator and multiple workers. The pseudocode is structured to reflect this division.
Coordinator As shown in Algorithm 5, the coordinator receives the graph size $n$, number of workers $M$, constant parameter $\psi>0$, privacy parameter $\varepsilon\in(0,1]$, privacy split fraction $f\in(0,1)$, and bias term $b$. It initializes three channels, $cRR,cMaxOut,cTCount$, to receive RR noisy edges, maximum noisy out-degrees, and local noisy triangle counts from workers (Algorithm 5). The coordinator manages the algorithm’s execution, collecting worker outputs and publishing updates each round. It first computes the low out-degree ordering, $Z$, using $k$-CoreD. In the first round, it launches $M$ asynchronous worker processes (Algorithm 5), each computing and sending noisy edges after RR. Before the second round, the coordinator aggregates and stores them in $X$, then publishes $X$ for global access (Algorithm 5), enabling workers to utilize the noisy public edges in subsequent computations. In the second round (Algorithm 5), workers compute noisy maximum out-degrees for their subgraphs using Algorithm 7. The coordinator then determines $\widetilde{d}_{\max}$, the maximum noisy out-degree across all workers (Algorithm 5). Finally, in the third round (Algorithm 5), each worker counts the triangles incident to its nodes using the low out-degree ordering, published noisy edges, and $\widetilde{d}_{\max}$. Workers send noisy local triangle counts to the coordinator, which aggregates them to compute the overall noisy triangle count (Algorithm 5).
[…]
Worker (Noisy Max Out-Degree) In Algorithm 7, workers maintain a variable $out_{max}$ which stores the maximum noisy out-degree of their subgraph. For each node $v$, the worker first computes the out-degree $d_{v}$ using the order provided in $Z$, where an edge $\left(v,j\right)$ is considered outgoing if $Z[j]>Z[v]$ (Algorithm 7). The out-degree $d_{v}$ is the number of outgoing edges from $v$. Then, the worker adds symmetric geometric noise with parameter $\varepsilon$ to $d_{v}$, computes the max noisy out-degree, and sends $out_{\max}$ to the coordinator (Algorithm 7).
Worker (Count Triangles) As shown in Algorithm 8, each worker computes the number of triangles incident to each node in its respective subgraph. For each node $v$, the outgoing edges are identified and sorted in ascending order by node IDs. The triangle count is determined by iterating over all unique pairs of outgoing neighbors $\{j,k\}$ of $v$, up to $\widetilde{d}_{\max}$ (Algorithm 8,8). For each pair, the triangle contribution is calculated as: $\frac{X_{\{j,k\}}\cdot(e^{\varepsilon}+1)-1}{e^{\varepsilon}-1}$, where $X_{\{j,k\}}$ represents the noisy presence (1) or absence (0) of an edge between $j$ and $k$ (Algorithm 8). To ensure privacy, additional noise is added to the triangle counts using the Laplace distribution [^6] We use Laplace noise here as it offers a smoother tradeoff for smaller parameters. with parameter $\frac{\varepsilon}{2\cdot\widetilde{d}_{\max}}$, where $\widetilde{d}_{\max}$ is the global maximum noisy out-degree. Upon completing the computation, each worker aggregates and returns the noisy triangle count for its entire subgraph (Algorithm 8).
Example 4.1.
In Fig. 4, we apply $\text{EdgeOrient}_{\Delta}$ to estimate the number of triangles incident to the blue node. The algorithm first orients the edges using a low out-degree ordering, so each node only considers neighbors with higher order. Randomized Response is then applied to the original adjacency list, and the resulting noisy edges are used in combination with the oriented edges to count triangles. As shown in the figure, while the blue node is part of two true triangles in the original graph, only one triangle is preserved under the noisy edges.
4.2. Theoretical Analysis
Memory Analysis & Communication Cost
Let $M$ be the number of workers and $n$ the graph size. Each worker processes $S$ nodes, where $S=\lfloor n/M\rfloor$ for $M-1$ workers, and the last worker handles $n-(M-1)\lfloor n/M\rfloor$ nodes. The coordinator manages three communication channels and publishes the noisy edges for the entire graph. The cRR structure, which aggregates noisy edges, requires $O(n^{2})$ space, while cTCount and cMaxOut, which collect triangle counts and maximum noisy out-degree, require $O(M)$ space each. Storing published noisy edges further adds $O(n^{2})$ space, resulting in a total coordinator memory requirement of $O(n^{2}+M)$. Each worker processes $O(S)$ nodes, requiring $O(Sn)$ space for the graph. The neighborsRR structure for storing noisy edges demands $O(Sn)$ space, while computing the maximum noisy out-degree requires $O(S)$. The final triangle count computation takes $O(S\cdot\widetilde{d}_{\max})$ space, where $\widetilde{d}_{\max}$ is the maximum noisy out-degree, leading to an overall worker memory requirement of $O(Sn)$. The algorithm runs three communication rounds beyond those for low out-degree ordering. Workers first send noisy edges, incurring $O(Sn)$ communication cost, followed by sending the maximum noisy out-degree and triangle counts, each requiring $O(M)$ communication. Thus, the total communication overhead for the algorithm is $O(n^{2}+M)$.
Privacy Guarantees
As before, our privacy guarantees are proven by implementing our triangle counting algorithm using local randomizers.
Lemma 4.2.
Our triangle counting algorithm is $\varepsilon$-LEDP.
Proof.
Our triangle counting algorithm calls Algorithm 1, which by Theorem 3.3 is $(\varepsilon/4)$-LEDP. Additionally, we release three sets of information, each of which we show to be $(\varepsilon/4)$-LEDP.
First, each node applies Randomized Response to the upper triangular adjacency matrix to generate a privacy-preserving set of edges. By (Dwork et al., 2006), this adjacency list output is a $(\varepsilon/4)$-local randomizer.
Second, each node releases its privacy-preserving out-degree. By Algorithm 7, the sensitivity of the out-degree (conditioning on $Z$) is 1 for neighboring adjacency lists. By the privacy of the geometric mechanism ((Balcer and Vadhan, 2018; Chan et al., 2011; Dwork et al., 2006, 2010a)), each node uses a $(\varepsilon/4)$-local randomizer to output its noisy degree.
Third, each node releases a privacy-preserving triangle count using its outgoing edges from the low out-degree ordering. To bound the sensitivity, we truncate the outgoing adjacency list (computed using $Z$) of each node by $\widetilde{d}_{\max}$. Given neighboring adjacency lists $\mathbf{a}$ and $\mathbf{a}^{\prime}$, assume $\mathbf{a}^{\prime}$ contains one additional neighbor $w$ (without loss of generality). Let $\overline{\mathbf{a}}$ and $\overline{\mathbf{a}^{\prime}}$ be the truncated adjacency lists. In the worst case, $\overline{\mathbf{a}}$ contains a node $u$ not in $\overline{\mathbf{a}^{\prime}}$, while $\overline{\mathbf{a}^{\prime}}$ contains $w$ (not in $\overline{\mathbf{a}}$). Let $j$ be defined as in Algorithm 8. If $j=u$, the first for-loop (Algorithm 8) counts at most $\widetilde{d}_{\max}$ additional triangles for $u$ (symmetrically for $w$). Assuming $u$ returns $\widetilde{d}_{\max}$ triangles and $w$ returns none, then for all other nodes $j\neq u$, the second for-loop (Algorithm 8) encounters at most $\widetilde{d}_{\max}$ additional triangles. Thus, the total difference in counted triangles between $\overline{\mathbf{a}}$ and $\overline{\mathbf{a}^{\prime}}$ is $2\widetilde{d}_{\max}$, giving a sensitivity of $2\widetilde{d}_{\max}$. By the privacy of the Laplace mechanism ((Dwork et al., 2006)), outputting local triangle counts is an $(\varepsilon/4)$-local randomizer.
Since the differing edge between neighboring graphs $G$ and $G^{\prime}$ affects at most one node’s out-degree, applying composition ((Dwork et al., 2006; Dwork and Lei, 2009; Dwork et al., 2010b)) over all four $(\varepsilon/4)$-local randomizers results in an $\varepsilon$-LEDP triangle counting algorithm. ∎
Approximation Guarantees
One of the major novelties in our proofs is via a new intricate use of the Law of Total Expectation and Law of Total Variance for the events where the out-degrees of each node is upper bounded by the noisy maximum out-degree $\widetilde{d}_{\max}$ (which is, in turn, upper bounded by the degeneracy $\widetilde{O}(d)$). Such use cases were unnecessarily in (Imola et al., 2022; Eden et al., 2023b) because they did not use oriented edges. We first upper bound the out-degree by $\widetilde{O}\left(\frac{d}{\varepsilon}\right)$.
Lemma 4.3.
Given a graph where edges are oriented according to Algorithm 8, the maximum out-degree of any node is at most $O\left(d+\frac{\log(D_{\max})\log^{2}(n)}{\varepsilon}\right)$.
Proof.
By Invariant 1 and Theorem 3.4 (in the supplementary materials), the out-degree of any node $v$ is at most $(2+\eta)k(v)+O\left(\frac{\log(D_{\max})\log^{2}(n)}{\varepsilon}\right)$, with high probability, where $k(v)$ is the core number of $v$. The largest core number is equal to the degeneracy of the graph (Seidman, 1983). Hence, the maximum out-degree of any node is upper bounded by $(2+\eta)d+O\left(\frac{\log(D_{\max})\log^{2}(n)}{\varepsilon}\right)=O\left(d+\frac{\log(D_{\max})\log^{2}(n)}{\varepsilon}\right)$
Lemma 4.4.
Given a graph where edges are oriented according to Algorithm 8, the number of oriented 4-cycles, denoted by $\overrightarrow{C}_{4}$, where each cycle contains two non-adjacent nodes with outgoing edges to the remaining two nodes (see Fig. 5), is at most $\widetilde{O}\left(\frac{n^{2}d^{2}}{\varepsilon^{2}}\right)$, with high probability, where $d$ is the degeneracy of the graph.
Proof.
There are $O(n^{2})$ unordered pairs of vertices $\{w,x\}$ that may serve as the black nodes in an oriented 4-cycle (see Fig. 5). For a fixed pair $\{w,x\}$, define $S_{w,x}=\{u\in V:w\to u\text{ and }x\to u\}$ to be the set of vertices that are the outgoing endpoints of the outgoing edges from both $w$ and $x$.
Under the low out-degree orientation computed by Algorithm 8 and by Lemma 4.3, each vertex has at most $O\left(d+\frac{\log(D_{\max})\log^{2}(n)}{\varepsilon}\right)$ out-neighbors and thus can have at most $O\left(\left(d+\frac{\log(D_{\max})\log^{2}(n)}{\varepsilon}\right)^{2}\right)=O\left(d^{2}+\frac{\log^{2}(D_{\max})\log^{4}(n)}{\varepsilon^{2}}\right)$ outgoing red pairs. Each pair $\{u,v\}\subseteq S_{w,x}$ forms an oriented 4-cycle with $\{w,x\}$, contributing at most this many cycles per black pair of vertices. Summing over all $O(n^{2})$ black vertex pairs yields a total of at most $O\left(n^{2}\left(d^{2}+\frac{\log^{2}(D_{\max})\log^{4}(n)}{\varepsilon^{2}}\right)\right)=\widetilde{O}\left(\frac{n^{2}d^{2}}{\varepsilon^{2}}\right)$ oriented 4-cycles. ∎
Figure 5. Oriented cycle of length $4$; two non-adjacent black nodes have edges oriented toward the remaining red nodes.
Lemma 4.5.
In expectation, our algorithm returns a $2$-approximation of the true triangle count: $\frac{(n^{3}-1)\cdot T}{n^{3}}\leq\mathbb{E}[\widetilde{\Delta}]\leq T$.
Proof.
We first prove that, in expectation, $\mathbb{E}\left[\frac{X_{j,k}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}\right]=\mathbbm{1}_{j,k}$ where $\mathbbm{1}_{j,k}$ is the indicator variable for whether edge $\{j,k\}$ exists in the original private graph. Since randomized response flips the bit indicating the existence of an edge between each pair of vertices with probability $\frac{1}{e^{\varepsilon/4}-1}$, we can simplify the expression to be:
$\mathbb{E}\left[\frac{X_{j,k}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}\right]$ $=\frac{\mathbb{E}[X_{j,k}]\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}.$
When $\mathbbm{1}_{j,k}=1$, the probability that we obtain a bit of $1$ is $\frac{e^{\varepsilon/4}}{e^{\varepsilon/4}+1}$ and our expression simplifies to
$\frac{\mathbb{E}[X_{j,k}]\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}$ $=\frac{\frac{e^{\varepsilon/4}}{e^{\varepsilon/4}+1}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}=1.$
When $\mathbbm{1}_{j,k}=0$, the probability that we obtain a bit of $1$ is $\frac{1}{e^{\varepsilon/4}+1}$ then our expression simplifies to
$\frac{\mathbb{E}[X_{j,k}]\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}$ $=\frac{\frac{1}{e^{\varepsilon/4}+1}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}=0.$
The expected value of the random variable obtained for each edge $e$ using our randomized response procedure is equal to $\mathbbm{1}_{e}$.
Now, we condition on the event that $\widetilde{d}_{\max}$ upper bounds the out-degree of every vertex after computing and conditioning on $D$ (the ordering). Let $U$ be this event.
Then, for each node, we compute all pairs of its outgoing neighbors and use the random variable indicating existence of the edge spanned by the endpoints to count every triangle composed of a pair of outgoing edges. Let $T_{v,j,k}$ be the random variable representing the existence of the queried triangle for node $v$ and outgoing edges $(v,j)$ and $(v,k)$. Then, $\mathbb{E}[T_{v,j,k}\mid U]=\mathbb{E}[X_{j,k}]$ since both outgoing edges $(v,j)$ and $(v,k)$ exists. By what we showed above, it then follows that $\mathbb{E}[T_{v,j,k}]=\mathbbm{1}_{j,k}$.
We must account for the symmetric geometric noise. Since the expectation of the symmetric geometric noise is 00 and we add together all of the values for each of the drawn noises, the expected total noise is 00 by linearity of expectations. Each triangle has a unique node that queries it since every triangle has a unique orientation of edges where two edges are outgoing from a vertex in the triangle. Hence, the expected sum of all queried triangles is $T$, conditioned on $U$, by linearity of expectations.
Finally, we remove the conditioning on the event $U$. Recall that $\mathbb{E}[W]=\sum_{y}\mathbb{E}[W\mid Y=y]\cdot\mathsf{Pr}(Y=y)$. Since truncation can only decrease the number of queried triangles (and hence the expectation), we upper and lower bound $0\leq\mathbb{E}[\widetilde{\Delta}\mid\neg U]\leq T$. Hence we only need to figure out the probability $P(U)$ to upper and lower bound $\mathbb{E}[\widetilde{\Delta}]$. The probability of event $U$ is the probability that $\max\left(\{d_{v}+\mathsf{Geom}(\varepsilon/4)\mid v\in V\}\right)+\frac{\log(n)}{\varepsilon}\geq\max(\{d_{v}\mid v\in V\})$ where $d_{v}$ is the out-degree of node $v$ given order $D$. This probability is lower bounded by the probability that $\max(\{d_{v}\mid v\in V\})+\mathsf{Geom}(\varepsilon/4)+\frac{c\log(n)}{\varepsilon}\geq\max(\{d_{v}\mid v\in V\})$ for a fixed constant $c\geq 1$; this means we want to lower bound $\mathsf{Pr}\left(\left(Q+\frac{c\log(n)}{\varepsilon}\right)\geq 0\right)$ where $Q\sim\mathsf{Geom}(\varepsilon/4)$. By concentration of random variables chosen from the symmetrical geometric distribution, we know that $\mathsf{Pr}\left(\left(Q+\frac{c\log(n)}{\varepsilon}\right)\geq 0\right)\geq 1-\frac{1}{n^{3}}$ for large enough constant $c\geq 1$. Specifically, setting $c=3$ gives us this bound. Hence, we can upper and lower bound $\frac{(n^{3}-1)\cdot T}{n^{3}}\leq\mathbb{E}[\widetilde{\Delta}]\leq T$. ∎
To calculate the variance of the triangle count obtained from our algorithm, we use a quantity denoted by $\overrightarrow{C}_{4}$ which is the number of oriented $4$-cycles where there exists two non-adjacent nodes with outgoing edges to the other two nodes in the cycle. See Fig. 5 for an example. The number of oriented cycles of length $4$ is upper bounded by $n^{2}d^{2}$ where $d$ is the degeneracy (maximum core number) in the graph, resulting in significant gains in utility over previous results which use the number of total (not oriented) $4$-cycles which could be as large as $\Omega(n^{4})$.
Lemma 4.6.
Our triangle counting algorithm returns a count with variance $O\left(\frac{nd^{2}\log^{6}n}{\varepsilon^{4}}+\overrightarrow{C}_{4}\right)$ where $d$ is the degeneracy (max core number) of the graph, $\overrightarrow{C}_{4}$ is the number of oriented cycles of length $4$, and $T$ is the number of (true) triangles in the private graph.
Proof.
As before, we first calculate the variance conditioned on the event $U$ that $\widetilde{d}_{\max}$ upper bounds the out-degrees of every node. Then, we use the law of total variance to remove this condition.
For notation simplicity, we omit the condition on $U$ from the right hand sides of the below equations. First, the variance of $\mathrm{Var}[X_{\{i,j\}}]=\frac{e^{\varepsilon}}{(e^{\varepsilon}+1)^{2}}$, since $X_{\{i,j\}}$ is a Bernoulli variable. Then, let $\hat{T}$ be our returned triangle count. The variance
$\mathrm{Var}\left[\hat{T}\mid U\right]$
$=\mathrm{Var}\left[\sum_{v\in[n]}\left(\sum_{j,k\in Out(v)}\left(\frac{X_{j,k}\cdot(e^{\varepsilon/4}+1)-1}{e^{\varepsilon/4}-1}\right)+\text{Lap}\left(\frac{\varepsilon}{2\widetilde{d}_{\max}}\right)\right)\right].$
Recall $\widetilde{d}_{\max}$ is our noisy maximum out-degree of any vertex. Then, our variance simplifies to
$\left(\frac{e^{\varepsilon/4}+1}{e^{\varepsilon/4}-1}\right)^{2}\cdot\mathrm{Var}\left[\sum_{v\in[n]}\sum_{j,k\in Out(v)}X_{j,k}\right]+\mathrm{Var}\left[\sum_{v\in[n]}\text{Lap}\left(\frac{\varepsilon}{2\widetilde{d}_{\max}}\right)\right].$
By the variance of the Laplace distribution, we can compute
$\mathrm{Var}\left[\sum_{v\in[n]}\text{Lap}\left(\frac{\varepsilon}{2\widetilde{d}_{\max}}\right)\right]\leq n\cdot\frac{8\widetilde{d}_{\max}^{2}}{\varepsilon^{2}}.$
Now, it remains to compute $\mathrm{Var}\left[\sum_{v\in[n]}\sum_{j,k\in Out(v)}X_{j,k}\right]$. The covariance is 00 if two queried pairs do not query the same $X_{j,k}$. The covariance is non-zero only in the case of queries which overlap in $X_{i,j}$. This occurs only in the case of an oriented $4$-cycle where $X_{i,j}$ is shared between two queries. Let $P_{2}$ be the set of all pairs of such queries that share $X_{j,k}$. In this case, the covariance is upper bounded by $\mathbb{E}[X_{j,k}^{2}]$. Hence, we can simplify our expression to be
(1) $\mathrm{Var}\left[\sum_{v\in[n]}\sum_{j,k\in Out(v)}X_{j,k}\right]$
(2) $\leq\sum_{v\in[n]}\sum_{j,k\in Out(v)}\mathrm{Var}\left[X_{j,k}\right]+2\cdot\sum_{T_{v,j,k},T_{w,j,k}\in P_{2}}\left(\mathbb{E}[X_{j,k}^{2}]\right)$
(3) $\leq n\cdot\widetilde{d}_{\max}^{2}\cdot\frac{e^{\varepsilon/4}}{(e^{\varepsilon/4}+1)^{2}}+\overrightarrow{C}_{4}\cdot\left(\frac{e^{\varepsilon/4}}{e^{\varepsilon/4}+1}\right),$
where $\overrightarrow{C}_{4}$ indicates the number of directed cycle of length $4$ (Fig. 5).
Hence, our total variance is upper bounded by
$\frac{8n\widetilde{d}_{\max}^{2}}{\varepsilon^{2}}+n\cdot\widetilde{d}_{\max}^{2}\cdot\frac{e^{\varepsilon/4}}{(e^{\varepsilon/4}+1)^{2}}+\overrightarrow{C}_{4}\cdot\frac{e^{\varepsilon/4}}{e^{\varepsilon/4}+1}.$
Finally, by the guarantees of our $k$-core decomposition algorithm, the maximum out-degree $d_{\max}$ is bounded by $d_{\max}\leq(2+\eta)\cdot d+O\left(\frac{\log(D_{\max})\log^{2}n}{\varepsilon}\right)$, with high probability, where $d$ is the degeneracy of the input graph and $D_{\max}\leq n$ is the maximum degree in the graph. Finally, we also know that the noise generated for $\widetilde{d}_{\max}$ is upper bounded by $O\left(\frac{\log n}{\varepsilon}\right)$. Thus, $\widetilde{d}_{\max}\leq(2+\eta)\cdot d+O\left(\frac{\log^{3}n}{\varepsilon}\right)$ Hence, our variance is upper bounded by $O\left(\frac{nd^{2}\log^{6}n}{\varepsilon^{4}}+\overrightarrow{C}_{4}\right)$.
We now remove our condition on $U$ and use the law of total variance to compute our unconditional variance. Recall the law of total variance states $\mathrm{Var}\left[Y\right]=\mathbb{E}[\mathrm{Var}\left[Y\mid X\right]]+\mathrm{Var}\left[\mathbb{E}[Y\mid X]\right]$. We computed $\mathrm{Var}\left[\hat{T}\mid U\right]$ above. We now compute $\mathrm{Var}\left[\hat{T}\mid\neg U\right]$. The main difference between when the event $U$ occurs and does not occur is that some adjacency lists of the outgoing neighbors will be truncated. Consequently, we sum over fewer $X_{j,k}$ variables in Eqs. 1 and 2. Thus, the variance when $U$ does not occur is upper bounded by the variance when $U$ does occur.
Now we calculate $\mathrm{Var}\left[\mathbb{E}[\hat{T}\mid U]\right]=\mathbb{E}[\mathbb{E}[\hat{T}\mid U]^{2}]-\mathbb{E}[\mathbb{E}[\hat{T}\mid U]]^{2}$. By our calculation in the proof of Lemma 4.5, we can calculate $\mathbb{E}[\hat{T}\mid U]=T$ and let $\frac{(n^{3}-1)\cdot T}{n^{3}}\leq Y=\mathbb{E}[\hat{T}\mid\neg U]\leq T$. Then,
$\mathbb{E}[\mathbb{E}[\hat{T}\mid U]^{2}]-\mathbb{E}[\mathbb{E}[\hat{T}\mid U]]^{2}$ $=\left(\frac{T^{2}}{2}+\frac{Y^{2}}{2}\right)-\left(\frac{T}{2}+\frac{Y}{2}\right)^{2}$
$=\frac{T^{2}+Y^{2}}{2}-\frac{T^{2}+2TY+Y^{2}}{4}$
$=\frac{T^{2}+Y^{2}-2TY}{4}$
$=\left(\frac{T-Y}{2}\right)^{2}$
$\leq\left(\frac{T-\frac{(n^{3}-1)\cdot T}{n^{3}}}{2}\right)^{2}$
$\leq\left(\frac{T}{2n^{3}}\right)^{2}$
$\leq\frac{1}{4}.$
Thus, our final variance is $O\left(\frac{nd^{2}\log^{6}n}{\varepsilon^{4}}+\overrightarrow{C}_{4}\right)$. ∎
Theorem 4.7.
With high constant probability, our triangle counting algorithm returns a $\left(1+\eta,O\left(\frac{\sqrt{n}d\log^{3}n}{\varepsilon^{2}}+\sqrt{\overrightarrow{C}_{4}}\right)\right)$-approximation of the true triangle count.
Proof.
We use Chebyshev’s inequality with the standard deviation calculated from Lemma 4.6. The $(1+\eta)$-approximation comes from our $\left(1-\frac{1}{n^{3}}\right)$-approximation of the expectation. ∎
See Table 2 for the $d$ values of real-world graphs; when $d=O(1)$ is constant, as is the case for real-world graphs, then $\overrightarrow{C}_{4}=O(n^{2})$ and $T=O(n)$. We improve previous theoretical additive errors from $O\left(\frac{\sqrt{C_{4}}}{\varepsilon}+\frac{n^{3/2}}{\varepsilon^{2}}\right)$ (Imola et al., 2022) to $O\left(\sqrt{\overrightarrow{C}_{4}}+\frac{\sqrt{n}\log^{3}n}{\varepsilon^{2}}\right)$, an improvement of at least a $\Omega(\sqrt{n})$ factor, translating to massive practical gains.
5. Experimental Evaluation
In this section, we evaluate the performance and accuracy of our $k$-core decomposition ($k$-CoreD) and triangle counting ($\text{EdgeOrient}_{\Delta}$) algorithms under Local Edge Differential Privacy (LEDP) using a distributed simulation. We benchmark against prior LEDP algorithms, and to highlight the limitations of Randomized Response (RR), we additionally implement RR-based baselines for both problems. All RR baselines are evaluated purely in terms of accuracy, as they are centralized algorithms and not directly comparable in runtime to our distributed setting. Furthermore, RR introduces significant computational overhead due to increased graph density from edge perturbation. Consequently, we omit RR results on many large graphs: $k$-core baselines fail due to out-of-memory (OOM) errors, while triangle counting exceeds timeout limits. These failures occur independently, underscoring the instability and inefficiency of naive RR methods on large-scale graphs.
$k$-Core Baselines. We compare $k$-CoreD against the LEDP $k$-core decomposition algorithm of (Dhulipala et al., 2022) (denoted $k$-Core), which we implement. Additionally, we construct an RR-based baseline (denoted $k$-CoreRR) that runs the standard peeling algorithm on the RR-perturbed graph and applies a scaling factor to correct the induced degrees, ensuring unbiased estimates. Our method achieves significantly better accuracy, reducing approximation error by up to two orders of magnitude over $k$-CoreRR and outperforming $k$-Core (Dhulipala et al., 2022) in both accuracy and efficiency—reducing the number of rounds by nearly two orders of magnitude across all graphs. $k$-CoreRR consistently fails on larger datasets due to memory exhaustion caused by increased graph density.
[…]
Table 2. Graph size, maximum core number, and number of triangles.
Graph Name Num. Vertices Num. Edges Max. Degree Max. Core Num. ($d$) Num. Triangles
email-eu-core 986 1,329,336 345 34 105,461
wiki 7115 100,761 1065 35 608,387
enron 36,692 183,830 1,383 43 727,044
brightkite 58,228 214,078 1,134 52 494,728
ego-twitter 81,306 1,342,296 3,383 96 13,082,506
gplus 107,614 12,238,285 20,127 752 1,073,677,742
stanford 281,903 1,992,635 38,625 71 11,329,473
dblp 317,080 1,049,866 343 113 2,224,385
brain 784,262 267,844,669 21,743 1200
orkut 3,072,441 117,185,083 33,313 253
livejournal 4,846,609 42,851,237 20,333 372
twitter 41,652,230 1,202,513,046 2,997,487 2488
friendster 65,608,366 1,806,067,135 5214 304
(a) Average Approximation Factor
(b) $80^{th}$ Percentile Approximation Factor
(c) $95^{th}$ Percentile Approximation Factor
Figure 6. $k$-core Decomposition Results.
Figure 7. $k$-Core Decomposition Avg. Response Time
Figure 8. $k$-Core Decomposition Number of Rounds
(a) gplus
(b) wiki
(c) livejournal
(d) gplus
(e) wiki
(f) livejournal
Figure 9. Avg. approximation factor for $k$-Core decomposition vs. epsilon ($\varepsilon$) and split fraction ($f$).
Experimental Setup. To evaluate our algorithms in a distributed simulation, we partition the input graph across $M$ worker processors and a single coordinator processor. Each worker handles a subset of nodes and their full adjacency lists, running LEDP algorithms locally. Workers communicate their privacy-preserving outputs to the coordinator, which aggregates the data and broadcasts new public information. This proceeds over multiple synchronous rounds, simulating a real-world distributed setting. We use $\mathbf{80}$ worker processors and a single coordinator. Each value is averaged over $\mathbf{5}$ runs, with a 4-hour wall-clock limit per run.
Parameters We use $\varepsilon=1.0$, bias term $=8$, approximation factor $(2+\eta)=5.625$ (matching non-private $k$-core experiments (Liu et al., 2022b)), and privacy split fraction $f=0.8$ (allocating $0.8\cdot\varepsilon$ to thresholding and $0.2\cdot\varepsilon$ to level moving step). We also conduct an ablation study on key parameters $\varepsilon$ and $f$, and our theoretical proofs show the approximation falls within a $(2+\eta)$-multiplicative factor.
Compute Resources We run experiments on a Google Cloud c3-standard-176 instance (3.3 GHz Intel Sapphire Rapids CPUs, 88 physical cores, 704 GiB RAM) with hyper-threading disabled. The code, implemented in Golang (Donovan and Kernighan, 2015), is publicly available (git, 2024).
Datasets We test our algorithms on a diverse set of 13 real-world undirected graphs from SNAP (Leskovec and Krevl, 2014), the DIMACS Shortest Paths Challenge road networks (Demetrescu et al., 2008), and the Network Repository (Rossi and Ahmed, 2015): email-eu-core, wiki, enron, brightkite, ego-twitter, gplus, stanford, dblp, orkut, livejournal, and friendster. We also use twitter, a symmetrized version of the Twitter network (Kwak et al., 2010), and brain, a highly dense human brain network from NeuroData (https://neurodata.io/). We remove duplicate edges, zero-degree vertices, and self-loops. Table 2 reflects the graph statistics after this removal. Exact triangle counts for some graphs are omitted due to time or memory constraints.
(a) Average Response Time
(b) Relative Error
(c) Average Approximation Factor
Figure 10. Triangle Counting Results.
(a) gplus
(b) stanford
(c) dblp
Figure 11. Avg. approx factor for triangle counting vs. epsilon ($\varepsilon$).
5.1. $k$-Core Decomposition
Response Time. Fig. 7 shows the response times of $k$-CoreD across all datasets. Our algorithm efficiently processes large-scale graphs, including billion-edge datasets like twitter and friendster, within four hours. These results validate the scalability and practicality of $k$-CoreD, demonstrating the impact of degree thresholding and bias terms in delivering both theoretical and practical improvements.
While measuring response time, a direct runtime comparison with the $k$-Core algorithm (Dhulipala et al., 2022) is not meaningful. That algorithm does not include bias correction, which often results in large negative noise causing most nodes to remain stuck at level 0. Since the algorithm proceeds for a fixed number of communication rounds (matching the levels in the LDS), but no nodes move beyond the first level, negligible work is performed in subsequent rounds—making the runtime unrealistically low while returning (the same) poor approximation for many nodes. Instead, we compare the number of communication rounds. As shown in Fig. 8, $k$-CoreD reduces the number of rounds by two orders of magnitude compared to the baseline, aligning with our theoretical bound of $O(\log(n)\cdot\log(D_{\max}))$. This translates to significantly lower communication overhead and improved scalability in distributed settings.
Accuracy We calculate the approximation factor for each node as $a_{v}=\frac{\max\left(s_{v},t_{v}\right)}{\min\left(s_{v},t_{v}\right)}$, where $s_{v}$ is the approximate core number and $t_{v}$ is the true core number. We use this metric to be consistent with the best-known non-private $k$-core decomposition implementations (Liu et al., 2022b; Dhulipala et al., 2017). These individual node approximation factors facilitate the computation of aggregate metrics: the average, maximum, $80^{th}$, and $95^{th}$ percentile approximation factors for each graph. The theoretical approximation bound in the absence of noise is calculated as $(2+\eta)$, which is $5.625$ for all graphs (labeled Theoretical Bound). Additionally, we adjust this bound to account for noise by incorporating the additive error term $\frac{\log_{1+\eta/5}^{3}\left(D_{max}\right)}{\varepsilon}$, where $n$ is the number of nodes, $D_{max}$ is the maximum degree (labeled Theoretical Bound (With Noise)). For this bound, we compute the effect of the additive noise on the multiplicative factor by adding $\frac{\log_{1+\eta/5}^{3}\left(D_{max}\right)}{\varepsilon\cdot k_{\max}}$, where $k_{\max}$ is the maximum core number, to $5.625$. Note that such a theoretical bound is a lower bound on the effect of the additive error on the multiplicative factor; for smaller core numbers, e.g. $k_{\min}<<k_{\max}$, the additive error leads to a much greater factor.
Figs. 6(a), 6(b) and 6(c) present the approximation factors achieved by our approach ($k$-CoreD) compared to the baseline $k$-core algorithm ($k$-Core) from (Dhulipala et al., 2022), and the randomized response baseline ($k$-CoreRR), along with theoretical bounds across various datasets. On average, our method maintains approximation factors below $\mathbf{4x}$ across all datasets, with the $80^{\text{th}}$ percentile staying under $\mathbf{5.5x}$, as illustrated in Fig. 6(a) and Fig. 6(b), demonstrating significantly lower variance compared to the baselines. These results remain well within the theoretical bounds without noise. Compared to $k$-Core (Dhulipala et al., 2022), $k$-CoreD consistently achieves better or comparable performance. Notably, for graphs such as brain and gplus, $k$-CoreD reduces the approximation factors from $131.55$ to $4.11$ and from $52.71$ to $3.27$, improvements of over 31x and 16x, respectively. Similarly, for orkut and wiki, our algorithm improves the approximation factors by 6.6x and 1.9x, respectively. Compared to $k$-CoreRR, our algorithm achieves consistently better performance across all datasets—often by two to three orders of magnitude. Notably, on the dblp graph, $k$-CoreD improves the approximation factor by over 47x. These results underscore the limitations of naive RR-based methods, which suffer from inflated graph density due to edge perturbation.
However, for many graphs, the difference between $k$-Core (Dhulipala et al., 2022) and our approach is less pronounced. This is due to the $k$-Core algorithm’s inability to move nodes up levels in the LDS, which results in an approximate core number of $2.5$ for most nodes. Given that real-world graphs often exhibit small core numbers (Table 2), the average approximation factor for $k$-Core becomes skewed, particularly for graphs with a significant proportion of low-core nodes. In contrast, for graphs with larger core numbers, such as gplus, brain, and orkut, our method demonstrates significant improvements. The advantages of our algorithm become more evident when examining the $80^{\text{th}}$ and $95^{\text{th}}$ percentile approximation factors. As shown in Fig. 6(b), our method consistently achieves a notable reduction in the $80^{\text{th}}$ percentile error compared to $k$-Core, with reductions of up to 42x, 40x, and 7.7x for graphs with large core numbers, such as brain, gplus, and orkut, respectively. Similarly, in the $95^{\text{th}}$ percentile error (Fig. 6(c)), our method achieves reductions of nearly 49x, 81x, and 7x for the same graphs. These results underscore the robustness and scalability of our algorithm in handling diverse graph structures with varying core number distributions.
Ablation Study We analyze the effect of varying the privacy parameter, $\varepsilon$, and the privacy split fraction, $f$, on the utility of the $k$-core decomposition algorithm by plotting the average approximation factor across different datasets: gplus, wiki, and livejournal, for our algorithm and the baseline. From the results shown in Fig. 9, we observe that the approximation factor improves as $\varepsilon$ increases, which aligns with the theoretical expectations of differential privacy where higher $\varepsilon$ allows for less noise and greater utility.
Further, we note that an optimal value of $0.8$ consistently minimizes the approximation factor across all datasets (Fig. 9). This is the case since degree thresholding affects the amount of noise that will be added per level for later computations. Thus, we use $f=0.8$ for all other experiments, as it strikes a balance between the two steps, ensuring better overall performance.
5.2. Triangle Counting
[…]
Accuracy Following the evaluation methodology in (Imola et al., 2022), we compute relative error for a graph using $\frac{\lvert\widetilde{\Delta}-\Delta\rvert}{\Delta}$, where $\widetilde{\Delta}$ represents the approximated triangle count and $\Delta$ the true triangle count. Additionally, we apply the theoretical bounds from Theorem 4.7 to our analysis. According to Fig. 10(b), our algorithm, $\text{EdgeOrient}_{\Delta}$, consistently achieves relative errors ranging from $\mathbf{10^{-1}}$ to $\mathbf{10^{-2}}$ across all datasets, remaining well within the theoretical bounds.
[…]
On small or dense graphs like wiki or email-eu-core, the TCountRR—which incurs a dominant error term growing proportional to $\frac{n^{3/2}}{\varepsilon^{2}}$—performs similarly to ours because $n$ is small (Eden et al., 2023a). However, in larger graphs, $n^{3/2}$ explodes, whereas our error depends instead on $\sqrt{n}\,d$ (with $d\ll n$). TCountRR must also consider all $O(n^{4})$ $4$-cycles, but our method only needs $\widetilde{O}(n^{2}d^{2})$ $4$-cycles. Consequently, on larger datasets such as enron and brightkite, we reduce the relative error (Fig. 10(b)) by roughly 17x and 45.5x, respectively. Compared to using 60 worker cores, we achieve up to a $1.41$x speedup.
To further analyze the limitations of prior approaches, we compute the multiplicative approximation factor of the triangle count, defined as $\frac{\max\left(\widetilde{\Delta},\Delta\right)}{\max\left(1,\min\left(\widetilde{\Delta},\Delta\right)\right)}$. Unlike relative error, this metric explicitly accounts for cases where algorithms, such as the one in (Imola et al., 2022), produce negative triangle counts, which severely undermines their utility in real-world scenarios. As shown in Fig. 10(c), our algorithm achieves consistently small approximation factors across all graphs, remaining within $\mathbf{[1.01,1.93]}$, and reducing the factor by six orders of magnitude compared to (Imola et al., 2022) and TCountRR. This highlights the robustness of our approach, $\text{EdgeOrient}_{\Delta}$, in maintaining low and stable approximation factors across diverse graph structures.
[…]
6. Conclusion
Large-scale network analysis often raises privacy concerns for sensitive data. We employ local edge differential privacy (LEDP), letting nodes protect their edges without a trusted authority. We propose novel LEDP algorithms for $k$-core decomposition and triangle counting that surpass prior work in accuracy and theoretical guarantees and introduce the first distributed framework to simulate these algorithms on a single machine. Experiments show our $k$-core approximations meet non-private theoretical bounds on average, while triangle counting errors are nearly two orders of magnitude lower than previous LEDP methods, with similar runtimes. Ongoing work extends this framework to multi-machine testing with real communication to capture added latency and explores multi-coordinator variants to eliminate single points of failure and reduce bottlenecks. These enhancements will support a broader range of LEDP algorithms. Our open-source framework (at (git, 2024)) invites the community to build on LEDP graph algorithms.